### v1

# Robust Permissive Controller Synthesis for Interval MDPs

Abstract
We address the problem of robust permissive controller synthesis for robots operating under uncertain dynamics, modeled as Interval Markov Decision Processes (IMDPs). IMDPs generalize standard MDPs by allowing transition probabilities to vary within intervals, capturing epistemic uncertainty from sensing noise, actuation imprecision, and coarse system abstractions—common in robotics. Traditional controller synthesis typically yields a single deterministic strategy, limiting adaptability. In contrast, permissive controllers (multi-strategies) allow multiple actions per state, enabling runtime flexibility and resilience. However, prior work on permissive controller synthesis generally assumes exact transition probabilities, which is unrealistic in many robotic applications. We present the first framework for robust permissive controller synthesis on IMDPs, guaranteeing that all strategies compliant with the synthesized multi-strategy satisfy reachability or reward-based specifications under all admissible transitions. We formulate the problem as mixed-integer linear programs (MILPs) and propose two encodings: a baseline vertex-enumeration method and a scalable dualization-based method that avoids explicit enumeration. Experiments on four benchmark domains show that both methods synthesize robust, maximally permissive controllers and scale to large IMDPs with up to hundreds of thousands of states.
I Introduction
Decision-making under uncertainty is a central challenge in robotics [1, 2]. Robots operate in environments where sensing is noisy, actuation is imprecise, and system abstractions are inevitably coarse. Standard Markov decision processes (MDPs) are widely used to model sequential decision-making, but they assume exact transition probabilities, which are rarely available in practice. Recent surveys highlight the broad landscape of uncertainty models [3, 4].
Among these models, Interval MDPs (IMDPs) play a critical role. By allowing transition probabilities to vary within intervals, IMDPs capture epistemic uncertainty while retaining a tractable structure for analysis and synthesis. Robust controller synthesis for IMDPs has been studied extensively, with applications ranging from robotic planning and verification to hybrid and neural network dynamics [5, 6, 7, 8]. However, existing methods typically return a single policy—often deterministic or randomized—but not permissive, which can be overly restrictive in practice.
In contrast, permissive controllers, also known as multi-strategies, allow multiple actions to remain enabled in each state. Such controllers preserve flexibility at runtime: a robot may adaptively select among admissible actions based on external conditions, unforeseen disturbances, or higher-level objectives, while still maintaining correctness. Permissive synthesis has been studied for MDPs and probabilistic games with exact transitions, often via mixed-integer linear programming (MILP) encodings that maximize permissiveness subject to property satisfaction [9, 10, 11, 12]. Yet, existing permissive methods assume precise probability models and do not account for epistemic uncertainty as captured by IMDPs.
The lack of robust permissive synthesis methods creates a significant gap for robotics applications. Consider a mobile robot navigating in an uncertain environment. A deterministic controller may commit to a single action sequence that is safe under all transition realizations, but in doing so, it may forgo many other safe alternatives. A permissive controller, in contrast, enables all actions consistent with safety, allowing adaptation at runtime. To be useful in robotics, such permissive strategies must be robust to transition uncertainty, guaranteeing correctness under all admissible distributions of the IMDP.
Another way to interpret permissive controllers is as shields for safe reinforcement learning (RL) [12, 13], which filter unsafe actions online while leaving other choices unrestricted. A permissive multi-strategy serves precisely this role, specifying admissible actions in each state and blocking the rest. Prior permissive synthesis methods thus align with shields in known MDPs, while robust IMDP synthesis aligns with worst-case reasoning under uncertainty. Bridging these ideas, robust permissive synthesis over IMDPs naturally yields robust shields: multi-strategies that guarantee safety against all admissible transitions while minimally restricting learning or heuristic policies. This view is particularly relevant for emerging paradigms such as meta-RL [14] and in-context RL [15], where agents must adapt across distributions of MDPs. Both are attracting interest in robotics, but ensuring safety under distributional variability remains challenging; robust permissive shields offer a principled way to address this while preserving flexibility for adaptation.
In this paper, we present the first framework for robust permissive controller synthesis on IMDPs. We focus on reachability and reward-based specifications, which are widely used in robotics. Our approach guarantees that every strategy compliant with the synthesized multi-strategy satisfies the specification under all admissible transitions. We formulate the synthesis problem as mixed-integer linear programs and propose two alternative encodings:
  1. 1.
    Vertex enumeration encoding, a direct formulation that is conceptually simple but introduces exponentially many constraints in the number of successors; and
  2. 2.
    Dualization-based encoding, which avoids explicit enumeration by leveraging robust optimization, yielding a compact formulation that scales linearly with the number of successors.
We implement both encodings in a prototype tool and evaluate them on four representative benchmark domains: aircraft collision avoidance, semi-autonomous vehicle control, obstacle navigation, and warehouse navigation. Our experiments show that both methods synthesize robust, maximally permissive controllers, scaling to IMDPs with hundreds of thousands of states.
By bridging permissive multi-strategy synthesis with robust IMDP reasoning, our work advances the state of the art in uncertainty-aware decision-making for robotics. It provides both theoretical foundations and practical algorithms for synthesizing controllers that are robust to model uncertainty while retaining runtime flexibility.
II Related Work
IMDPs in Robotics. Interval MDPs have been widely applied as abstractions in robotic planning and control under uncertainty. Recent work has used IMDPs to model dynamics with learned uncertainty for navigation and control tasks, including legged locomotion [5], safe motion planning [6], and switched stochastic systems [16]. IMDP abstractions have also been constructed from perception pipelines, yielding conservative bounds with probabilistic guarantees [17]. Other works demonstrate the use of IMDPs in stochastic hybrid systems [7], neural-network dynamics [8], and compositional abstractions for scalability [18]. Data-driven approaches extend this further, with IMDP abstractions learned from samples and equipped with PAC guarantees [19, 20]. These studies highlight the versatility of IMDP abstractions for robotics, though they primarily focus on model construction and verification rather than permissive controller synthesis.
IMDP Controller Synthesis. Robust controller synthesis for IMDPs has been extensively studied, with algorithms addressing temporal-logic objectives, reachability, and multi-objective properties. Classical results established robust dynamic programming under interval uncertainty [21, 22], while subsequent work developed robust synthesis for probabilistic temporal logics [23, 24]. Recent research has focused on scalability, richer objectives and learning, including multi-objective synthesis [25], PAC-based learning of IMDPs [26] and efficient PAC-bounded abstractions for dynamical systems [27]. Compositional and structural approaches enable scalable synthesis for high-dimensional systems [18], and tool support continues to advance [28, 29, 30]. While these works solve robust synthesis for IMDPs under various objectives, they all yield deterministic or randomized strategies, limiting runtime flexibility.
Permissive Controller Synthesis. Permissive controller synthesis was introduced to preserve flexibility by allowing multiple actions per state while still satisfying specifications. Initial work encoded permissive multi-strategy synthesis as a MILP with penalties for disabled actions [9]. Subsequent studies extended this idea to multi-objective settings, synthesizing permissive controllers that accommodate uncertain human preferences [11], and to robust planning in models with set-valued transitions under temporal-logic objectives [31]. Beyond formal synthesis, permissive controllers are closely related to shields in safe reinforcement learning, which block unsafe actions online while leaving other choices unrestricted; in this setting, a permissive multi-strategy naturally serves as a shield that enforces safety with minimal interference [12, 13]. While these approaches demonstrate the utility of permissive strategies, they assume exact or set-valued probabilities. To the best of our knowledge, no prior work addresses robust permissive synthesis for IMDPs under interval uncertainty, which is the gap filled by this paper.
III Problem Formulation
Interval MDPs (IMDPs). Formally, an IMDP is a tuple $\mathcal{M}=(S,s_{\iota},A,\check{P},\hat{P},R)$, where $S$ is a finite set of states with initial state $s_{\iota}\in S$, $A$ is a finite set of actions, $\check{P},\hat{P}:S\times A\times S\to[0,1]$ assign lower and upper bounds to each transition, and $R:S\times A\to\mathbb{R}_{\geq 0}$ is a reward function. Intuitively, an IMDP represents a family of MDPs. For each state-action pair $(s,a)$, the admissible successor distributions are $\mathcal{P}(s,a)=\{\,P\in\Delta(S)\mid\check{P}(s,a,s^{\prime})\leq P(s,a,s^{\prime})\leq\hat{P}(s,a,s^{\prime})\;\;\forall s^{\prime}\in S\,\}.$ We define $\mathcal{P}$ as the set of all transition functions $P:S\times A\to\Delta(S)$ with $P(s,a,\cdot)\in\mathcal{P}(s,a)$ for every $(s,a)$.
Permissive Controller. An IMDP is controlled by a strategy (also called a policy), which in this work is assumed to be deterministic and memoryless. Formally, a strategy is a function $\sigma:S\to A$ such that $\sigma(s)\in\alpha(s)$, where $\alpha(s)\subseteq A$ denotes the set of enabled actions in state $s$. A permissive controller, or multi-strategy, generalizes this notion by allowing a set of actions at each state; formally, it is a mapping $\theta:S\to 2^{A}$ such that $\theta(s)\subseteq\alpha(s)$ and $\theta(s)\neq\emptyset$ for all $s\in S$. For a multi-strategy $\theta$, define its permissiveness as $\beta(\theta)\coloneqq\sum_{s\in S}\sum_{a\in\alpha(s)}\mathbf{1}[a\in\theta(s)]$. A strategy $\sigma$ is said to be compliant with a multi-strategy $\theta$, denoted $\sigma\triangleleft\theta$, if $\sigma(s)\in\theta(s)$ for every $s\in S$. For two multi-strategies $\theta$ and $\theta^{\prime}$, we write $\theta\sqsubset\theta^{\prime}$ if $\theta(s)\subset\theta^{\prime}(s)$ for all $s\in S$.
[…]
Satisfaction Semantics. Given an IMDP $\mathcal{M}$, a property $\varphi$, a strategy $\sigma$, and a transition function $P\in\mathcal{P}$, we write $(\mathcal{M},\sigma,P)\models\varphi$ if $\varphi$ holds in the MDP obtained by resolving all choices in $\mathcal{M}$ using strategy $\sigma$ and transition function $P$. For a multi-strategy $\theta$, $(\mathcal{M},\theta,P)\models\varphi$ if this holds for all $\sigma\triangleleft\theta$. Robust satisfaction then quantifies over all admissible transitions: $(\mathcal{M},\sigma)\models_{\text{rob}}\varphi$ if $(\mathcal{M},\sigma,P)\models\varphi$ for all $P\in\mathcal{P}$, and $(\mathcal{M},\theta)\models_{\text{rob}}\varphi$ if $(\mathcal{M},\theta,P)\models\varphi$ for all $P\in\mathcal{P}$.
Problem Statement. We aim to synthesize a robust permissive controller that guarantees satisfaction of the specification under all admissible uncertainties, while retaining as many action choices as possible.
Problem: Robust Permissive Controller Synthesis Given an IMDP $\mathcal{M}=(S,s_{\iota},A,\check{P},\hat{P},R)$ and a specification $\varphi$ of the form $\mathsf{P}_{\bowtie p}(\Diamond T)$ or $\mathsf{R}_{\bowtie b}(\Diamond T)$, synthesize a multi-strategy $\theta$ such that: 1. (Robust satisfaction) $(\mathcal{M},\theta)\models_{\text{rob}}\varphi$. 2. (Maximal permissiveness) There exists no multi-strategy $\theta^{\prime}$ with $\theta\sqsubset\theta^{\prime}$ such that $(\mathcal{M},\theta^{\prime})\models_{\text{rob}}\varphi$.
Example 1. Consider a robotic navigation task modeled as the IMDP $\mathcal{M}$ in Figure 1. A robot starts in $s_{0}$ and must reach the goal state $s_{3}$. From $s_{0}$, it may choose a fast action ($f$), which can reach $s_{3}$ directly but with risk of failure into $s_{2}$, or a medium action ($m$), which is safer but must be applied twice through $s_{1}$. Transition probabilities are specified as intervals of radius $\varepsilon\in[0,1]$, where $\varepsilon$ is a fixed constant capturing the level of epistemic uncertainty. For example, $P(s_{0},f,s_{3})\in[0.78-\varepsilon,\,0.78+\varepsilon]$, clipped to $[0,1]$. A property such as $\varphi=\mathsf{P}_{\geq 0.65}(\Diamond s_{3})$ asks whether the probability of eventually reaching the goal is at least $0.65$. The synthesis problem then asks for a multi-strategy $\theta$ such that $(\mathcal{M},\theta)\models_{\text{rob}}\varphi$ while being maximally permissive. $\blacksquare$
Figure 1: A simple robotic navigation task modeled as an IMDP.
IV Methods
We now present methods for solving the IMDP robust permissive controller synthesis problem. A straightforward approach is to explicitly enumerate all vertices of the uncertainty polytopes in the IMDP and adopt existing permissive controller synthesis techniques for MDPs [9]. This yields a mixed-integer linear program (MILP) encoding, but the number of vertices grows exponentially with the number of successor states per state–action pair. To address this scalability issue, we develop a dualization-based encoding that captures the adversarial choice of transition probabilities without explicit enumeration, yielding a compact MILP. We present the vertex enumeration encoding in Section IV-A and the dualization-based encoding in Section IV-B.
IV-A Vertex Enumeration Encoding
We present an MILP for the probabilistic reachability property $\mathsf{P}{\geq p}(\Diamond T)$, given in Encoding IV-A. The formulation introduces real-valued variables $x_{s}\in[0,1]$, denoting the minimal probability of reaching $T$ from state $s$, and binary variables $y_{s,a}$, indicating whether the multi-strategy $\theta$ admits action $a\in\alpha(s)$.
Encoding 1: Vertex Enumeration for $\mathsf{P}_{\geq p}(\Diamond T)$ $\operatorname*{maximize}\sum_{s\in S}\sum_{a\in\alpha(s)}y_{s,a}$ (1a) subject to: $\forall s\in S:\ \sum_{a\in\alpha(s)}y_{s,a}\geq 1$ (1b) $x_{s_{\iota}}\geq p$ (1c) $\forall s\in T:\ x_{s}=1$ (1d) $\forall s\in S\setminus T,\ \forall a\in\alpha(s),\ \forall v\in\mathcal{V}(s,a):$ $x_{s}\leq\sum_{s^{\prime}\in S}P^{(v)}(s,a,s^{\prime})\cdot x_{s^{\prime}}+M(1-y_{s,a})$ (1e)
The objective function (1a) maximizes the number of admitted actions, thereby ensuring maximal permissiveness. Constraint (1b) ensures that all states with enabled actions admit at least one of them. Constraint (1c) enforces that the initial state achieves the threshold $p$. Constraint (1d) fixes the value of all target states to 1.
Constraint (1e) enforces robustness by requiring $x_{s}$ to hold for all feasible successor distributions of each non-target pair $(s,a)$. The term $M(1-y_{s,a})$ deactivates the constraint when action $a$ is not allowed, where $M$ is a sufficiently large constant (the standard big-$M$ method) used for linearization and constraint deactivation. The vertex set $\mathcal{V}(s,a)$ consists of the extreme points of $\mathcal{P}(s,a)$, obtained by pushing transition probabilities to their interval bounds while maintaining that they remain nonnegative and sum to one. Each $v\in\mathcal{V}(s,a)$ defines a concrete distribution $P^{(v)}(s,a,\cdot)$, and enforcing the inequality for all $v$ ensures that $x_{s}$ is robust.
Example 2. We apply Encoding IV-A to the IMDP in Figure 1 with respect to the property $\varphi=\mathsf{P}_{\geq 0.65}(\Diamond s_{3})$. Variables $x_{s}$ record the minimal reachability probability from each state, while binary variables $y_{s_{0},f}$, $y_{s_{0},m}$, and $y_{s_{1},m}$ specify which actions (fast or medium) are allowed by the multi-strategy in states $s_{0}$ and $s_{1}$. The objective is to maximize the sum $y_{s_{0},f}+y_{s_{0},m}+y_{s_{1},m}$. Constraints (1b)–(1d) enforce that at least one action is chosen in $s_{0}$ and $s_{1}$, that $x_{s_{0}}\geq 0.65$, and that $x_{s_{3}}=1$. Constraint (1e) is instantiated for each vertex of the uncertainty polytope. For example, for the fast action at $s_{0}$, $\mathcal{P}(s_{0},f)$ reduces to an interval over successors $\{s_{2},s_{3}\}$ and yields two inequalities:
$x_{s_{0}}$ $\leq(0.22+\varepsilon)\,x_{s_{2}}+(0.78-\varepsilon)\,x_{s_{3}}+M(1-y_{s_{0},f}),$
$x_{s_{0}}$ $\leq(0.22-\varepsilon)\,x_{s_{2}}+(0.78+\varepsilon)\,x_{s_{3}}+M(1-y_{s_{0},f}).$
Solving the MILP shows how the resulting multi-strategy depends on the uncertainty radius $\varepsilon$ and the threshold $p$. For $\varepsilon=0.1$, requiring $\mathsf{P}_{\geq 0.65}(\Diamond s_{3})$ admits only the fast action. Lowering the threshold to $p=0.6$ or reducing uncertainty to $\varepsilon=0.05$ admits both fast and medium actions. These cases are illustrated in Figure 2, which plots the minimal (dashed) and maximal (solid) reachability probabilities for each action as the uncertainty radius $\varepsilon$ varies. $\blacksquare$
Figure 2: Robust reachability probability as the uncertainty radius $\varepsilon$ varies.
Correctness. The correctness of the MILP in Encoding IV-A is stated in the following theorem, with the proof provided in the appendix.
Theorem 1
Let $\mathcal{M}$ be an IMDP and $\varphi=\mathsf{P}_{\geq p}(\Diamond T)$ a property. There exists a robust, maximally permissive multi-strategy $\theta$ in $\mathcal{M}$ for $\varphi$ if and only if there exists an optimal assignment to the MILP in Encoding IV-A whose objective value equals the permissiveness of $\theta$, i.e., $\beta(\theta)=\sum_{s\in S}\sum_{a\in\alpha(s)}y_{s,a}$.
Adapting Encoding IV-A for Other Specifications. For properties of the form $\mathsf{P}_{\leq p}(\Diamond T)$, constraint (1c) is replaced with $x_{s_{\iota}}\leq p$, and constraint (1e) becomes
$x_{s}\geq\sum_{s^{\prime}\in S}P^{(v)}(s,a,s^{\prime})\cdot x_{s^{\prime}}-M(1-y_{s,a}).$
For expected total reward properties $\mathsf{R}_{\geq b}(\Diamond T)$, we instead use $x_{s}\in\mathbb{R}_{\geq 0}$ to denote reward-to-go variables, modify constraint (1c) to $x_{s_{\iota}}\geq b$, set constraint (1d) to $x_{s}=0$, and replace constraint (1e) with
$x_{s}\leq\sum_{s^{\prime}\in S}P^{(v)}(s,a,s^{\prime})\cdot x_{s^{\prime}}+r(s,a)+M(1-y_{s,a}),$
where $r(s,a)$ is the immediate reward of action $a$ in state $s$. Analogous changes apply to properties of the form $\mathsf{R}_{\leq b}(\Diamond T)$ by reversing the inequalities accordingly. We also impose constraints on set $T$ being reached with probability 1.
Scalability. Although conceptually straightforward, the vertex enumeration encoding can scale poorly. The number of vertices $|\mathcal{V}(s,a)|$ of each polytope $\mathcal{P}(s,a)$ can be exponential in the number of successor states, and constraint (1e) must be instantiated for every vertex. As a result, the MILP may contain exponentially many constraints in the size of the IMDP, making this encoding impractical for larger models and motivating the dualization-based approach as follows.
IV-B Dualization-Based Encoding
Encoding 2: Dualization for $\mathsf{P}_{\geq p}(\Diamond T)$ $\operatorname*{maximize}\ \sum_{s\in S}\sum_{a\in\alpha(s)}y_{s,a}$ (2a) subject to: $\forall s\in S:\ \sum_{a\in\alpha(s)}y_{s,a}\geq 1$ (2b) $x_{s_{\iota}}\geq p$ (2c) $\forall s\in T:\ x_{s}=1$ (2d) $\forall(s,a,s^{\prime}):\ \lambda_{s,a}-\hat{u}_{s,a,s^{\prime}}+\check{u}_{s,a,s^{\prime}}\leq x_{s^{\prime}}$ (2e) $\forall(s,a):\ x_{s}\leq M(1-y_{s,a})+\eta_{s,a}$ $+\sum_{s^{\prime}\in S}\bigl(\check{P}(s,a,s^{\prime})\,\check{u}_{s,a,s^{\prime}}-\hat{P}(s,a,s^{\prime})\,\hat{u}_{s,a,s^{\prime}}\bigr)$ (2f) $\forall(s,a):\ -M\,y_{s,a}\leq\eta_{s,a}\leq M\,y_{s,a}$ (2g) $\forall(s,a):\ \lambda_{s,a}-M(1-y_{s,a})\leq\eta_{s,a}$ $\leq\lambda_{s,a}+M(1-y_{s,a})$ (2h)
We now present a compact MILP encoding based on linear programming duality as shown in Encoding IV-B. The objective (2a) and constraints (2b)–(2d) are identical to those in Encoding IV-A.
Instead of enumerating all vertices of the uncertainty polytopes, we encode the worst-case choice of transition probabilities directly using dual variables, a standard approach in robust optimization [32]. Here, $\hat{u}_{s,a,s^{\prime}},\check{u}_{s,a,s^{\prime}}\geq 0$ are dual variables associated with the upper and lower bounds $\hat{P}(s,a,s^{\prime})$ and $\check{P}(s,a,s^{\prime})$, $\lambda_{s,a}$ is a free dual variable, and $\eta_{s,a}$ is an auxiliary variable used to linearize the product $\lambda_{s,a}y_{s,a}$ with the big-$M$ method [33].
Constraint (2e) enforces dual feasibility: for each state–action pair $(s,a)$ and successor $s^{\prime}$, the dual variables satisfy $\lambda_{s,a}-\hat{u}_{s,a,s^{\prime}}+\check{u}_{s,a,s^{\prime}}\leq x_{s^{\prime}}$. Constraint (2f) is the core robust inequality, bounding the value $x_{s}$ by the dual objective. Intuitively, it ensures that the worst-case distribution consistent with the intervals is captured without explicit vertex enumeration. Finally, constraints (2g)–(2h) implement the big-$M$ linearization, guaranteeing $\eta_{s,a}=0$ when $y_{s,a}=0$ and $\eta_{s,a}=\lambda_{s,a}$ when $y_{s,a}=1$.
Example 3. We revisit the IMDP and property from Example IV-A, now using Encoding IV-B. For the fast action at $s_{0}$, constraint (2e) introduces dual feasibility inequalities linking $\lambda_{s_{0},f},\hat{u}_{s_{0},f,s^{\prime}},\check{u}_{s_{0},f,s^{\prime}}$ to the successor values $x_{s^{\prime}}$. Constraint (2f) then yields the robust bound
$x_{s_{0}}\leq\;$ $M(1-y_{s_{0},f})+\eta_{s_{0},f}$
$+(0.22-\varepsilon)\,\check{u}_{s_{0},f,s_{2}}-(0.22+\varepsilon)\,\hat{u}_{s_{0},f,s_{2}}$
$+(0.78-\varepsilon)\,\check{u}_{s_{0},f,s_{3}}-(0.78+\varepsilon)\,\hat{u}_{s_{0},f,s_{3}}.$
Here the auxiliary variable $\eta_{s_{0},f}$ linearizes the product $\lambda_{s_{0},f}y_{s_{0},f}$, ensuring that $\lambda_{s_{0},f}$ contributes only when $y_{s_{0},f}=1$. Solving this encoding yields the same multi-strategy outcomes as shown in Figure 2. $\blacksquare$
Correctness. The correctness of the MILP in Encoding IV-B is stated in the following theorem, with the proof provided in the appendix.
Theorem 2
Let $\mathcal{M}$ be an IMDP and $\varphi=\mathsf{P}_{\geq p}(\Diamond T)$ a property. There exists a robust, maximally permissive multi-strategy $\theta$ in $\mathcal{M}$ for $\varphi$ if and only if there exists an optimal assignment to the MILP in Encoding IV-B whose objective value equals the permissiveness of $\theta$, i.e., $\beta(\theta)=\sum_{s\in S}\sum_{a\in\alpha(s)}y_{s,a}$.
Adapting Encoding IV-B for Other Specifications. For $\mathsf{P}_{\leq p}(\Diamond T)$, constraint (2c) is replaced by $x_{s_{\iota}}\leq p$, and constraint (2f) becomes
$x_{s}$ $\geq-M(1-y_{s,a})+\eta_{s,a}$
$+\sum_{s^{\prime}\in S}\bigl(\check{P}(s,a,s^{\prime})\,\check{u}_{s,a,s^{\prime}}-\hat{P}(s,a,s^{\prime})\,\hat{u}_{s,a,s^{\prime}}\bigr).$
For reward properties $\mathsf{R}_{\bowtie b}(\Diamond T)$, the variables $x_{s}$ represent reward-to-go, constraint (2c) is replaced by $x_{s_{\iota}}\bowtie b$, constraint (2d) is set to $x_{s}=0$ for all $s\in T$, and the immediate reward $r(s,a)$ is added in constraint (2f). In both cases, all other constraints remain unchanged.
Scalability. The dualization-based encoding may introduce more constraints in small instances such as the IMDP in Figure 1. However, while vertex enumeration grows exponentially with the number of successor states, dualization requires only a linear number. This advantage becomes evident in larger IMDPs, as demonstrated in Section V.
V Experiments
We have developed a prototype implementation of the proposed synthesis methods. The implementation leverages the PRISM model checker [28] to parse IMDP models and specifications, and the Gurobi Optimizer [34] to solve the resulting MILP encodings. All experiments were conducted on a laptop equipped with a 2.8 GHz Quad-Core Intel Core i7 processor and 16 GB of RAM.
V-A Benchmark Domains
We evaluated our methods on four representative IMDP benchmark domains, each capturing a different robotic decision-making scenario under uncertainty. In the first three domains, the specification is a probabilistic reachability property of the form $P_{\geq p}(\Diamond T)$, requiring that the target set $T$ is reached with probability at least $p$ under all admissible transitions. The final domain considers a reward-based specification of the form $R_{\leq b}(\Diamond T)$.
Aircraft Collision Avoidance (ACA). Adapted from [35], this domain models a controlled aircraft navigating a discretized airspace with an adversarial aircraft. The objective is to reach a target region while avoiding collisions with probability at least $p$ under all admissible uncertainties. We consider three IMDP instances, where the parameter sets the maximum number of successor states per state–action pair, controlling the branching factor.
Semi-Autonomous Vehicle (SAV). Adapted from [10], this domain models a ground vehicle navigating a grid world while maintaining intermittent communication with a controller. Communication occurs via two lossy channels, with message-loss probabilities depending on the relay location. At each step, the vehicle may either attempt communication (with limited retries) or move to an adjacent cell. The objective is to reach the opposite corner with probability at least $p$; failure occurs if the vehicle moves too far without successful intermediate communication. The model parameter controls the size of the grid world.
Obstacle Navigation (OBS). Inspired by the FrozenLake problem [36], this domain models an agent moving through a stochastic grid world containing frozen cells that permanently trap the agent. The agent must reach a goal region with probability at least $p$. The model parameters control both the grid dimension and the number of steps required for a move, thereby varying the granularity of the state space.
Warehouse Navigation (WH). Adapted from [11], this domain models a robot traversing a warehouse partitioned into zones connected by checkpoints. Unlike the previous domains, the specification is reward-based, requiring that the expected number of steps to reach the target zone is at most $b$ under all admissible uncertainties. The parameter specifies the number of steps needed to move between checkpoints.
Domain IMDP Size Vertex Enumeration (Encoding IV-A) Dualization (Encoding IV-B)
Name Parameters #States #Trans #Binary #Real #Constraints Time(s) Permissive(%) #Binary #Real #Constraints Time(s) Permissive(%)
ACA 10 328 4,739 984 390 110,398 4.3 98.6 [90.1] 984 11,774 31,010 7.7 98.6 [87.0]
12 453 6,498 1,203 529 585,749 22.7 98.9 [92.5] 1,203 15,855 41,507 5.6 98.9 [88.9]
14 640 9,583 1,632 733 2,814,507 145.4 99.4 [95.8] 1,632 23,070 60,163 4.3 99.4 [93.9]
SAV 5 279 999 727 363 1,370 0.6 97.0 [33.3] 727 3,731 8,274 5.3 97.0 [35.3]
6 411 1,503 1,099 539 2,050 46.3 97.6 [35.0] 1,099 5,615 15,755 218.6 97.6 [35.0]
7 567 2,103 1,543 747 2,858 2,564.7 97.6 [31.5] 1,543 7,859 22,071 4,030.2 97.6 [35.1]
OBS 5; 1000 73,025 73,299 73,083 73,030 146,471 2.1 99.9 [63.2] 73,083 365,789 951,108 4.8 99.9 [51.9]
6; 100 11,336 1,760 11,424 11,342 23,324 1.4 99.7 [42.3] 11,424 57,704 115,840 3.7 99.7 [31.8]
6; 1000 113,036 113,460 113,124 113,042 226,724 3.2 99.9 [50.0] 113,124 566,204 1,132,840 14.8 99.9 [67.4]
WH 250 2,762 3,785 3,768 2,762 6,549 0.2 99.9 [99.9] 3,768 17,868 36,861 1.1 99.9 [99.9]
500 5,512 7,535 7,518 5,512 13,049 0.4 99.9 [99.9] 7,518 35,618 63,261 1.8 99.9 [99.9]
1000 11,012 15,035 15,018 11,012 26,049 0.7 99.9 [99.9] 15,018 71,118 146,261 3.0 99.9 [99.9]
TABLE I: Experimental results comparing vertex enumeration (Encoding 1) and dualization (Encoding 2). Permissive (%) values are reported as normalized permissiveness, with the choice-state permissiveness shown in brackets.
V-B Experimental Setup
We compare the vertex enumeration (Encoding IV-A) and dualization-based (Encoding IV-B) MILP formulations across the benchmark domains. Table I reports the IMDP model size (number of states and transitions), MILP encoding size (number of binary variables, real-valued variables, and constraints), and solution time for each instance.
Recall that permissiveness $\beta(\theta)$ counts the total number of state–action pairs allowed by a multi-strategy. For reporting, we normalize this measure to obtain values in $[0,1]$. We use two variants: (i) normalized permissiveness, given by $\tfrac{\sum_{s\in S}|\theta(s)|}{\sum_{s\in S}|\alpha(s)|}$ over all states $S$, and (ii) normalized choice-state permissiveness, the same ratio restricted to non-target states that are reachable from the initial state and have at least two available actions ($|\alpha(s)|\geq 2$). The latter focuses on states where the controller faces genuine non-deterministic choices, offering a clearer view of the flexibility preserved by the synthesized multi-strategy.
V-C Results Analysis
The results in Table I show that both encodings are capable of synthesizing maximally permissive robust strategies. In all benchmark instances, the normalized permissiveness is close to 1, confirming that the synthesized strategies preserve nearly all available actions.
Both approaches also scale to very large IMDPs with up to hundreds of thousands of states and transitions, solving all but some SAV instances within seconds. SAV-7 instances are notably slower, due to their more complex structure and communication-related modeling overhead.
Regarding MILP size, both encodings introduce the same number of binary variables, but differ in continuous variables and constraints. The vertex enumeration method is conceptually straightforward and yields relatively compact encodings for small models. However, its number of constraints grows exponentially with the number of successors, resulting in very large MILPs and long runtimes on larger instances (e.g., ACA models). In contrast, the dualization approach scales linearly with the number of successors, leading to more balanced MILPs and shorter runtimes in such cases.
Finally, while both methods achieve the same normalized permissiveness across all benchmarks, their choice-state permissiveness often differs. This indicates that the two encodings can produce different multi-strategies, even though both satisfy the same robust permissiveness objective.
VI Conclusion
We presented the first framework for robust permissive controller synthesis on IMDPs, enabling controllers that guarantee specification satisfaction under all admissible transitions while preserving multiple action choices for runtime flexibility. The problem was formulated as mixed-integer linear programs with two encodings: a conceptually simple vertex-enumeration method and a dualization-based method that scales linearly with the number of successors. Experiments on four benchmark domains showed that both approaches synthesize maximally permissive robust strategies, with the dualization method offering superior scalability on large models.
Future work includes applying our methods to a wide range of robotic domains, and integrating permissive synthesis with learning-based control. In particular, robust permissive strategies align naturally with the notion of shielding in safe reinforcement learning, suggesting opportunities for safety-preserving integration with meta- and in-context RL, where agents must adapt across distributions of tasks.
Appendix
Correctness of Encoding IV-A. We first establish two auxiliary lemmas and then present the proof of Theorem 1.
Lemma 1
For any $(s,a)$ and vector $x\in\mathbb{R}^{S}$, the extremum of $\sum_{s^{\prime}}P(s,a,s^{\prime})\,x_{s^{\prime}}$ over $P\in\mathcal{P}(s,a)$ is attained at some $P^{(v)}(s,a,\cdot)\in\mathcal{V}(s,a)$.
Proof:
For fixed $(s,a)$ and $x$, the expression $\sum_{s^{\prime}}P(s,a,s^{\prime})\,x_{s^{\prime}}$ is linear, and $\mathcal{P}(s,a)$ is a convex polytope. Hence any extremum over $\mathcal{P}(s,a)$ is attained at a vertex $P^{(v)}(s,a,\cdot)\in\mathcal{V}(s,a)$. ∎
Lemma 2
Let $\mathcal{M}$ be an IMDP, $\varphi=\mathsf{P}_{\geq p}(\Diamond T)$ a property, and $\theta$ a multi-strategy. Consider the inequalities for $s\in S$: $x_{s}\leq\min_{a\in\theta(s)}\sum_{s^{\prime}\in S}P^{(v)}(s,a,s^{\prime})\cdot x_{s^{\prime}}$. Then values $\bar{x}_{s}\in\mathbb{R}$ for $s\in S$ are a solution to the above inequalities iff $\bar{x}_{s}=\inf_{\sigma\triangleleft\theta}{\mathit{Pr}}^{\sigma,v}_{\mathcal{M},s}(\Diamond T)$, where ${\mathit{Pr}}^{\sigma,v}_{\mathcal{M},s}(\Diamond T)$ denotes the probability of reaching the target set $T$ under strategy $\sigma$ and transition distribution $P^{(v)}$ of $\mathcal{M}$, starting from state $s$.
Proof:
Fixing a vertex $v$ yields a standard MDP with transitions $P^{(v)}$. The inequalities correspond to the Bellman equations for minimum reachability; by monotonicity and standard MDP results [37], their solution $\bar{x}_{s}$ coincides with the value function. ∎
Proof of Theorem 1: (Soundness) Let $(x,y)$ be feasible for Encoding IV-A and define $\theta(s)=\{a\mid y_{s,a}=1\}$. Constraints (1b)–(1d) enforce non-deadlock, $x_{s_{\iota}}\geq p$, and $x_{s}=1$ for $s\in T$. Constraint (1e), together with Lemmas 1 and 2, ensures that $x$ is a sub-fixed-point of the robust Bellman operator under $\theta$. Since $\bar{x}$ is the least fixed point of this operator, we obtain $x\leq\bar{x}$, hence $\bar{x}_{s_{\iota}}\geq x_{s_{\iota}}\geq p$. Optimality of the MILP objective yields maximal permissiveness.
(Completeness) Conversely, given a robust maximally permissive $\theta$, let $y$ encode $\theta$ and set $x=\bar{x}$, the robust reachability values from Lemma 2. Then $(x,y)$ satisfies all constraints, the objective equals $\beta(\theta)$, and optimality follows from maximal permissiveness. $\blacksquare$
Correctness of Encoding IV-B. We first establish an auxiliary lemma and then present the proof of Theorem 2.
Lemma 3
Fix $(s,a)$ and $x\in\mathbb{R}^{S}$. The robust inequality $x_{s}\;\leq\;\min_{P(\cdot\mid s,a)\in\mathcal{P}(s,a)}\sum_{s^{\prime}\in S}P(s,a,s^{\prime})\,x_{s^{\prime}}$ holds iff there exist dual variables $\hat{u}_{s,a,s^{\prime}},\check{u}_{s,a,s^{\prime}}\geq 0$ and $\lambda_{s,a}\in\mathbb{R}$ such that $\lambda_{s,a}-\hat{u}_{s,a,s^{\prime}}+\check{u}_{s,a,s^{\prime}}\;\leq\;x_{s^{\prime}}\quad(\forall s^{\prime}\in S)$ and $x_{s}\;\leq\;\lambda_{s,a}+\sum_{s^{\prime}\in S}\bigl(\check{P}(s,a,s^{\prime})\,\check{u}_{s,a,s^{\prime}}-\hat{P}(s,a,s^{\prime})\,\hat{u}_{s,a,s^{\prime}}\bigr).$
Proof:
The right-hand side of the robust inequality is the optimum of an LP over $P(s,a,\cdot)$ subject to interval and simplex constraints. Its dual has variables $\hat{u},\check{u}\geq 0$ and $\lambda$, yielding exactly the displayed conditions. By strong LP duality [32], the primal minimum equals the dual maximum, establishing equivalence. ∎
Proof of Theorem 2: (Soundness) Let $(x,y,\hat{u},\check{u},\lambda,\eta)$ be feasible for Encoding IV-B and set $\theta(s)=\{a\mid y_{s,a}=1\}$. Constraints (2b)–(2d) enforce non-deadlock, $x_{s_{\iota}}\!\geq\!p$, and $x_{s}\!=\!1$ for $s\!\in\!T$. By Lemma 3, (2e)–(2f) are equivalent to the robust Bellman inequality $x_{s}\leq\min_{P\in\mathcal{P}(s,a)}\sum_{s^{\prime}}P(s,a,s^{\prime})x_{s^{\prime}}$ for each allowed $(s,a)$; (2g)–(2h) linearize $\eta_{s,a}=\lambda_{s,a}y_{s,a}$ (big-$M$), so the dual bound is active only if $y_{s,a}=1$. Hence $x$ is a sub-fixed-point of the robust Bellman operator under $\theta$. Let $\bar{x}$ be the robust reachability values; then $x\leq\bar{x}$ by Lemma 2, so $\bar{x}_{s_{\iota}}\geq p$. Optimality of the objective (2a) yields maximal permissiveness.
(Completeness) Let $\theta$ be a robust maximally permissive multi-strategy. Set $y$ to encode $\theta$ and $x=\bar{x}$ (the robust values from Lemma 2). For each allowed $(s,a)$, Lemma 3 provides dual variables $\hat{u},\check{u},\lambda$ making (2e)–(2f) hold at $x$; take $\eta_{s,a}=\lambda_{s,a}$ if $y_{s,a}=1$ and $\eta_{s,a}=0$ otherwise, so (2g)–(2h) hold. Thus all constraints are satisfied and the objective equals $\beta(\theta)$; optimality follows from maximal permissiveness. $\blacksquare$

### v2

@@
-Robust Permissive Controller Synthesis for Interval MDPs
+Optimization-Based Robust Permissive Synthesis for Interval MDPs
 Abstract
-We address the problem of robust permissive controller synthesis for robots operating under uncertain dynamics, modeled as Interval Markov Decision Processes (IMDPs).
-IMDPs generalize standard MDPs by allowing transition probabilities to vary within intervals, capturing epistemic uncertainty from sensing noise, actuation imprecision, and coarse system abstractions—common in robotics.
-Traditional controller synthesis typically yields a single deterministic strategy, limiting adaptability.
-In contrast, permissive controllers (multi-strategies) allow multiple actions per state, enabling runtime flexibility and resilience.
-However, prior work on permissive controller synthesis generally assumes exact transition probabilities, which is unrealistic in many robotic applications.
-We present the first framework for robust permissive controller synthesis on IMDPs, guaranteeing that all strategies compliant with the synthesized multi-strategy satisfy reachability or reward-based specifications under all admissible transitions.
-We formulate the problem as mixed-integer linear programs (MILPs) and propose two encodings: a baseline vertex-enumeration method and a scalable dualization-based method that avoids explicit enumeration.
-Experiments on four benchmark domains show that both methods synthesize robust, maximally permissive controllers and scale to large IMDPs with up to hundreds of thousands of states.
+We present an optimization-based framework for robust permissive synthesis for Interval Markov Decision Processes (IMDPs), motivated by robotic decision-making under transition uncertainty.
+In many robotic systems, model inaccuracies and sensing noise lead to interval-valued transition probabilities.
+While robust IMDP synthesis typically yields a single policy and permissive synthesis assumes exact models, we show that robust permissive synthesis under interval uncertainty can be cast as a global mixed-integer linear program (MILP) that directly encodes robust Bellman constraints.
+The formulation maximizes a quantitative permissiveness metric (the number of enabled state–action pairs), while guaranteeing that every compliant strategy satisfies probabilistic reachability or expected reward specifications under all admissible transition realizations.
+To address the exponential complexity of vertex-based uncertainty representations, we derive a dualization-based encoding that eliminates explicit vertex enumeration and scales linearly with the number of successors.
+Experimental evaluation on four representative robotic benchmark domains demonstrates scalability to IMDPs with hundreds of thousands of states.
+The proposed framework provides a practical and general foundation for uncertainty-aware, flexibility-preserving controller synthesis in robotic systems.
 I Introduction
-Decision-making under uncertainty is a central challenge in robotics [1, 2].
-Robots operate in environments where sensing is noisy, actuation is imprecise, and system abstractions are inevitably coarse.
-Standard Markov decision processes (MDPs) are widely used to model sequential decision-making, but they assume exact transition probabilities, which are rarely available in practice.
-Recent surveys highlight the broad landscape of uncertainty models [3, 4].
-Among these models, Interval MDPs (IMDPs) play a critical role.
-By allowing transition probabilities to vary within intervals, IMDPs capture epistemic uncertainty while retaining a tractable structure for analysis and synthesis.
-Robust controller synthesis for IMDPs has been studied extensively, with applications ranging from robotic planning and verification to hybrid and neural network dynamics [5, 6, 7, 8].
-However, existing methods typically return a single policy—often deterministic or randomized—but not permissive, which can be overly restrictive in practice.
-In contrast, permissive controllers, also known as multi-strategies, allow multiple actions to remain enabled in each state.
-Such controllers preserve flexibility at runtime: a robot may adaptively select among admissible actions based on external conditions, unforeseen disturbances, or higher-level objectives, while still maintaining correctness.
-Permissive synthesis has been studied for MDPs and probabilistic games with exact transitions, often via mixed-integer linear programming (MILP) encodings that maximize permissiveness subject to property satisfaction [9, 10, 11, 12].
-Yet, existing permissive methods assume precise probability models and do not account for epistemic uncertainty as captured by IMDPs.
-The lack of robust permissive synthesis methods creates a significant gap for robotics applications.
-Consider a mobile robot navigating in an uncertain environment.
-A deterministic controller may commit to a single action sequence that is safe under all transition realizations, but in doing so, it may forgo many other safe alternatives.
-A permissive controller, in contrast, enables all actions consistent with safety, allowing adaptation at runtime.
-To be useful in robotics, such permissive strategies must be robust to transition uncertainty, guaranteeing correctness under all admissible distributions of the IMDP.
-Another way to interpret permissive controllers is as shields for safe reinforcement learning (RL) [12, 13], which filter unsafe actions online while leaving other choices unrestricted.
-A permissive multi-strategy serves precisely this role, specifying admissible actions in each state and blocking the rest.
-Prior permissive synthesis methods thus align with shields in known MDPs, while robust IMDP synthesis aligns with worst-case reasoning under uncertainty.
-Bridging these ideas, robust permissive synthesis over IMDPs naturally yields robust shields: multi-strategies that guarantee safety against all admissible transitions while minimally restricting learning or heuristic policies.
-This view is particularly relevant for emerging paradigms such as meta-RL [14] and in-context RL [15], where agents must adapt across distributions of MDPs.
-Both are attracting interest in robotics, but ensuring safety under distributional variability remains challenging; robust permissive shields offer a principled way to address this while preserving flexibility for adaptation.
-In this paper, we present the first framework for robust permissive controller synthesis on IMDPs.
-We focus on reachability and reward-based specifications, which are widely used in robotics.
-Our approach guarantees that every strategy compliant with the synthesized multi-strategy satisfies the specification under all admissible transitions.
-We formulate the synthesis problem as mixed-integer linear programs and propose two alternative encodings:
+Decision-making under uncertainty is a fundamental challenge in robotics [35, 20].
+Robots operate under sensing noise, actuation imprecision, and modeling errors, yet standard Markov decision processes (MDPs) assume exact transition probabilities.
+Among uncertainty models, Interval MDPs (IMDPs) capture epistemic uncertainty by allowing transition probabilities to vary within intervals, while remaining amenable to formal analysis and synthesis [4, 33].
+Robust controller synthesis for IMDPs has been widely studied in robotic planning, verification, and learning-based control [17, 18, 8, 1].
+However, existing methods typically synthesize a single policy—deterministic or randomized—which can be overly restrictive in practice.
+In contrast, permissive controllers (multi-strategies) retain multiple admissible actions per state, preserving flexibility at runtime while maintaining correctness.
+Permissive synthesis has been developed for MDPs and probabilistic games with exact transition models, often via mixed-integer linear programming formulations that maximize permissiveness subject to specification satisfaction [11, 19, 9, 16].
+Yet these approaches assume precise transition probabilities and do not address epistemic uncertainty as captured by IMDPs.
+For robotics applications, permissiveness must therefore be combined with robustness, ensuring correctness under all admissible transition realizations.
+Permissive controllers can also be viewed as shields for safe reinforcement learning (RL) [16, 2], filtering unsafe actions online while leaving other choices unrestricted.
+Recent work has used IMDP abstractions to construct data-driven shields for unknown continuous systems [30], highlighting the promise of interval-based models for safety-critical learning.
+Robust permissive synthesis over IMDPs naturally unifies these ideas, yielding shields that provide worst-case safety guarantees while minimally restricting adaptive or learning-based policies.
+This perspective is particularly relevant for emerging paradigms such as meta-RL [5] and in-context RL [25], where agents must operate across varying environments while maintaining safety.
+In this paper, we present a framework for robust permissive controller synthesis on IMDPs.
+We consider probabilistic reachability and expected reward specifications, which are widely used in robotics.
+Our approach guarantees that every strategy compliant with the synthesized multi-strategy satisfies the specification under all admissible transition probabilities.
+We formulate the synthesis problem as mixed-integer linear programs and develop two alternative encodings:
   1. 1.
@@
   2. 2.
-    Dualization-based encoding, which avoids explicit enumeration by leveraging robust optimization, yielding a compact formulation that scales linearly with the number of successors.
-We implement both encodings in a prototype tool and evaluate them on four representative benchmark domains: aircraft collision avoidance, semi-autonomous vehicle control, obstacle navigation, and warehouse navigation.
-Our experiments show that both methods synthesize robust, maximally permissive controllers, scaling to IMDPs with hundreds of thousands of states.
-By bridging permissive multi-strategy synthesis with robust IMDP reasoning, our work advances the state of the art in uncertainty-aware decision-making for robotics.
-It provides both theoretical foundations and practical algorithms for synthesizing controllers that are robust to model uncertainty while retaining runtime flexibility.
+    Dualization-based encoding, which avoids explicit vertex enumeration via robust optimization, yielding a compact formulation that scales linearly with the number of successors.
+We implement both encodings in a prototype tool and evaluate them on four benchmark domains: aircraft collision avoidance, semi-autonomous vehicle control, obstacle navigation, and warehouse navigation.
+The results demonstrate that both approaches synthesize robust and highly permissive controllers, scaling to IMDPs with hundreds of thousands of states.
+By unifying permissive multi-strategy synthesis with robust IMDP reasoning, this work advances uncertainty-aware decision-making for robotics.
+It provides both theoretical guarantees and practical algorithms for synthesizing controllers that remain correct under model uncertainty while preserving runtime flexibility.
 II Related Work
@@
 Beyond formal synthesis, permissive controllers are closely related to shields in safe reinforcement learning, which block unsafe actions online while leaving other choices unrestricted; in this setting, a permissive multi-strategy naturally serves as a shield that enforces safety with minimal interference [16, 2].
-While these approaches demonstrate the utility of permissive strategies, they assume exact or set-valued probabilities.
-To the best of our knowledge, no prior work addresses robust permissive synthesis for IMDPs under interval uncertainty, which is the gap filled by this paper.
+These approaches, however, assume exact or set-valued transition models and do not address interval uncertainty.
+Recent work has extended shielding ideas to IMDP abstractions of unknown continuous systems [30], where unsafe actions are iteratively pruned using value iteration until a worst-case violation probability threshold is satisfied.
+The resulting shield is inclusion-maximal with respect to the safety bound and focuses on safe LTL specifications.
+In contrast, we formulate robust permissive synthesis for IMDPs as a single global optimization problem.
+Our approach encodes robust Bellman constraints directly in a mixed-integer linear program and maximizes a permissiveness metric (the total number of enabled state–action pairs), while supporting probabilistic reachability and expected reward objectives.
 III Problem Formulation
@@
 OBS 5; 1000 73,025 73,299 73,083 73,030 146,471 2.1 99.9 [63.2] 73,083 365,789 951,108 4.8 99.9 [51.9]
-6; 100 11,336 1,760 11,424 11,342 23,324 1.4 99.7 [42.3] 11,424 57,704 115,840 3.7 99.7 [31.8]
+6; 100 11,336 11,760 11,424 11,342 23,324 1.4 99.7 [42.3] 11,424 57,704 115,840 3.7 99.7 [31.8]
 6; 1000 113,036 113,460 113,124 113,042 226,724 3.2 99.9 [50.0] 113,124 566,204 1,132,840 14.8 99.9 [67.4]
@@
 V-B Experimental Setup
+Our evaluation focuses on assessing the scalability and optimality of the proposed MILP encodings; direct comparison with IMDP-based shielding approaches such as [30] is outside our scope due to differences in problem formulation and objectives.
 We compare the vertex enumeration (Encoding IV-A) and dualization-based (Encoding IV-B) MILP formulations across the benchmark domains.
@@
 V-C Results Analysis
-The results in Table I show that both encodings are capable of synthesizing maximally permissive robust strategies.
+The results in Table I show that both encodings successfully synthesize maximally permissive robust strategies.
 In all benchmark instances, the normalized permissiveness is close to 1, confirming that the synthesized strategies preserve nearly all available actions.
@@
 VI Conclusion
-We presented the first framework for robust permissive controller synthesis on IMDPs, enabling controllers that guarantee specification satisfaction under all admissible transitions while preserving multiple action choices for runtime flexibility.
+We presented a framework for robust permissive controller synthesis on IMDPs, enabling controllers that guarantee specification satisfaction under all admissible transitions while preserving multiple action choices for runtime flexibility.
 The problem was formulated as mixed-integer linear programs with two encodings: a conceptually simple vertex-enumeration method and a dualization-based method that scales linearly with the number of successors.